# thanks: Current address: Laboratoire de Physique de l’École normale supérieure, Université PSL, CNRS, Sorbonne Université and Université Paris Cité, 75005 Paris, France

### v1

@@
We investigate the wetting transitions displayed by the collection of active Brownian particles (ABPs) confined within rigid, impenetrable, flat walls. In our computational study using Brownian dynamics simulations, the wall-particle interactions are implemented with a short-range repulsive potential. An enhanced rotational diffusion at the walls is used as a control parameter for wetting transitions. Increasing the wall rotational diffusion destabilizes complete wetting, and the aggregate shows morphological transitions. We observe a sequence of morphological transitions with an increase in wall rotational diffusion: symmetric complete wetting (SCW), asymmetric complete wetting (ACW), partial wetting (PW) with droplet formation, and drying. In the PW state, we compute the contact angle as a function of activity and rotational noise. Our analysis indicates that these transitions are linked to enhanced kinetic energy fluctuations of particles and bubble formations in the dense state. We further characterize the nature of these transitions by systematically analyzing an order parameter. Our work shows that modifying local reorientation rates alone is sufficient to induce wetting transitions in active systems.
@@ I Introduction
While the properties of passive fluids on solid substrates have been extensively studied, there is a growing interest in exploring the aggregation and spreading of several active and biological systems on solid substrates and membranes Douezan_PNAS_2011 ; Perez-Gonzalez_NatPhys_2019 ; pallares2023stiffness ; Kusumaatmaja_JCB2021 ; Agudo-Canalejo2021 ; Alert_Langmuir_2019 ; Zhao_2021 ; melaugh2016shaping ; ponisch2019bacterial ; fins2024steer . In addition, the spread of bacterial species on substrates has also been studied in this context melaugh2016shaping ; ponisch2019bacterial . Unlike their passive counterparts, such living systems continuously convert energy into mechanical motion and hence far from equilibrium. The collective behavior of such self-motile systems has been theoretically studied using minimal particle-based models. It is well known that such systems display bulk phase separation at sufficiently large density and motility, called Motility-induced phase separation (MIPS) Redner2013 ; Fily2012 ; Cates2015 . They also display wall accumulation under confinement, even in the absence of explicit adhesive interactions Bechinger2016 ; lee2017interface ; Elgeti2013 . Several numerical studies have investigated the nature of wall aggregation of self-propelling particles sepulveda2017wetting ; SepulvedaSoto_PRE2018_activewetting ; neta2021wetting ; Turci_PRL_2021 ; Das2020_SM ; Das2020_PRE ; Turci_SM_2024 . It has been shown that the particle aggregates display a continuous transition from a complete wetting to a partial wetting state on rigid, impenetrable walls as the particle reorientation rate is increased sepulveda2017wetting ; SepulvedaSoto_PRE2018_activewetting . Such wetting transitions of active aggregates are also observed when the substrate properties are altered. The effect of wall permeablilty Turci_PRL_2021 ; Turci_SM_2024 and porosity Das2020_SM ; Das2020_PRE are explored in some detail. An intriguing morphological transition is observed in the aggregates of particles on the porous walls. With a slight increase in the particle crossing rate, the aggregates display a morphological transition from a uniformly spread state to a liquid droplet with a non-zero contact angle Das2020_SM ; Das2020_PRE , displaying a striking similarity to equilibrium wetting transitions. Apart from systems of self-propelling particles, studies have also been conducted on wetting transitions in other non-equilibrium models Hinrichsen_PRL_1997 ; Hinrichsen_PRE_2003 .
@@ II Model and simulation details
We have conducted the simulations with system sizes varying from $N=4455$ to $N=38400$, while keeping the area fraction $\varphi=N\pi\sigma^{2}/4L_{x}L_{y}$ fixed at $\varphi\simeq 0.3$. At this density, the probability of bulk phase separation is negligible and the clustering occurs only at the walls. Most of our simulations are conducted for $N=38400$ in a rectangular box with $L_{x}/\sigma=617$ and $L_{y}/\sigma=162$. The simulations were run up to time $T/t_{0}=3100$, for a range of $Pe=0-300$, maximum up to $10$ independent runs. For simulations, we use a time step $10^{-5}t_{0}$ for $Pe<180$ and $5\times 10^{-6}t_{0}$ for $Pe\geq 180$. We vary the rotational diffusion near walls $D_{r}^{w}$ in the range of $0-4335$. The upper limit of $D_{r}^{w}$ is limited such that the time-step of integration is at least an order lesser than the time to rotate for escaping a flat surface $\pi/2D_{r}^{w}$. We observe a variety of morphological states as we vary $D_{r}^{w}$ for different Pe.
@@ III Results
Figure 1: (a) Examples of cluster morphology with rigid, impenetrable walls for $N=38400$ and $Pe=240$: (i) symmetric complete wetting ($D_{r}^{w}=30$), (ii) asymmetric complete wetting ($D_{r}^{w}=135$), (iii) partial wetting ($D_{r}^{w}=375$) (b) Two-dimensional steady-state density distribution, with a color bar denoting the local density $\rho(x,y)$ of the states in (a). (c) Vacancy fraction $\ell$ as a function of $D_{r}^{w}$ for five independent runs. The error bars denote the standard deviation over time. The symbols denote symmetric complete wetting (SCW)($\blacksquare$), asymmetric complete wetting (ACW)(), partial wetting (PW)(
@@
When $D_{r}^{w}$is increased, keeping the Pe same, particles no longer aggregate symmetrically on both walls; instead, we observe a uniform aggregation on one of the walls and no aggregation on the opposite wall (see Fig. 1(a)(ii)). In this state, one of the liquid-vapor interfaces formed in SCW is replaced by a wall-vapor interface. Although a simultaneous occurrence of all three interfaces (liquid-vapor, wall-liquid, and wall-vapor) is a feature of a partial wetting state, the wet wall is covered by a uniform, dense layer of particles, unlike what was typically observed in equilibrium partial wetting (see supplementary MOVIE2). We also observe that the thickness of such a dense layer is approximately twice that of the aggregate thickness formed in SCW and the fraction of particles in the dense state is approximately the same. We term this state, asymmetric complete wetting (ACW). Such asymmetric states have previously been observed in simulations of Lennard-Jones fluids nijmeijer1990wetting ; sikkenk1988simulation
@@ III.1 Density profile and vacancy fraction
Although the instantaneous steady-state configurations clearly indicate a morphological transition of the dense states, we also calculate the time-averaged density distribution $\rho(x,y)$ of ABPs for the characterization of these states (see Appendix B for details). As shown in Fig. 1(b)(i)-(iii), $\rho$ provides a clear distinction between liquid and vapor states and relatively smooth liquid-vapor interfaces for all the morphological states. While SCW and ACW states provide a flat liquid-vapor interface, the PW state shows a curved interface.
We further calculate the vacancy fraction $\ell$, which quantifies the fraction of the length of the wall where the dense layer of particles is absent. In other words, $\ell$ quantifies the fractional length of the wall-vapor interfaces (see Appendix B for details). The value of $\ell$ is an indicator of the type of morphological state of the dense aggregates. If $\ell\simeq 0$, the system is in SCW state. In the case of ACW, where one of the walls is dry, $\ell\simeq 0.5$. For partial wetting (droplets) $10.5<\ell<$ 1 and for completely dry state $\ell\simeq 1$.
@@ III.2 State diagram
The morphological states of the dense aggregates shown in Fig. 1 are for a fixed activity ($Pe=240$). However, the qualitative behavior of the system is also influenced by the particle activity. In Fig. 2, we summarize the states as a function of $Pe$ and $D_{r}^{w}$. A novel feature observed in the simulation is the formation of the droplet in the PW state, making it crucial to identify the parameter range where partial wetting occurs. Therefore, we introduce a slightly different definition for vacancy fraction, called the wet-wall vacancy fraction $\ell_{w}$, which is calculated only for walls that are partially or completely wet. With this definition, both SCW and ACW are treated as morphologically equivalent states with the same $\ell_{w}$. Thus, ($\ell_{w}\simeq 0$) for both SCW and ACW and for PW state, $\ell_{w}>0$. However, we keep track of ACW and SCW by estimating $\ell$ and summarize this information in Fig. 2.
Figure 2: State diagram showing steady-state morphology of dense clusters with the variation of activity $Pe$ and enhanced rotational diffusion near the wall $D_{r}^{w}$ for the number of active particles, $N=38400$. The color coding in the panel indicates the wet-wall vacancy fraction, $\ell_{w}$ considered for the wet wall cases, except for states showing complete drying for all realizations. The system exhibits symmetric complete wetting $\blacksquare$ (Fig.1(a)(i)), asymmetric complete wetting (Fig.1(a)(ii)), partial wetting (Fig.1(a)(iii))
@@
For $Pe\lesssim 60$, we do not observe any aggregation, and the system is in a completely dry state for all $D_{r}^{w}$. For higher activity ($Pe\gtrsim 100$), we observe SCW states for low $D_{r}^{w}$($D_{r}^{w}$$\lesssim 100$) and other aggregated states at higher $D_{r}^{w}$. The droplet formations in PW states are observed for a range of $D_{r}^{w}$ for sufficiently large Pe, which is distinguished by a higher $\ell_{w}$. This range becomes narrower as $Pe$ is decreased until it disappears for $Pe\lesssim 100$. This region is also marked by the existence of bi-stable states, where PW states coexist with ACW states. In the state diagram Fig. 2, we observe the appearance of re-entrant wetting behavior for $D_{r}^{w}\geq 1500$.
@@ III.3 Contact angle of droplets
Once we identify the parameter range where the droplet states are observed, we analyze how the droplet morphology is influenced by both $Pe$ and $D_{r}^{w}$. For this, we use the time-averaged density profile, as it clearly identifies the liquid-vapor interface and its shape. First, we note that for a fixed $D_{r}^{w}$, increasing activity leads to a change in droplet shape, as illustrated in Fig. 3(a). For small $Pe$, the droplet is relatively flat upon increasing the $Pe$, and the droplet curvature increases until it saturates at large $Pe$ (Fig.3(a)). We also calculate the contact angle $\theta$ from the local tangent of the liquid-vapor interface at the liquid-wall contact point for different parameters. The contact angle is low for a relatively flat aggregate, and it increases as the interfacial curvature increases (see Fig. A4). In Fig. 3(b), we plot $\theta$ as a function of $D_{r}^{w}$ for a fixed $Pe$. Similar to the trend observed with the change in $Pe$, we find that $\theta$ is sensitive to an increase in $D_{r}^{w}$ for relatively small values of $D_{r}^{w}$, and it saturates around $\pi/6$ at large $D_{r}^{w}$. We note that we never observe $\theta$ approaching $\pi/2$ for any of the simulation parameters.
Figure 3: (a) Time-averaged density profile for droplets keeping $D_{r}^{w}=540$ for increasing $Pe$, showing gradual increase in curvature of droplets and the contact angle. (b) Contact angle $\theta$ (in °$\degree$) for droplets for increasing $D_{r}^{w}$ for $Pe=240$ and $Pe=300$. The dotted lines are guide to the eye.
@@ III.4 Analysis of asymmetric wetting transition
We have observed from the simulations that the dense state often encompasses ‘bubbles’, temporarily formed compact domains with low density (see Fig. A5). Qualitative analyses of simulation results indicate an increase in the number of bubbles formed within the dense layer. Such bubble formations are reported inside the dense regions formed by MIPS and in the presence of walls  martinroca2021mips ; caporusso2020mips ; shi2020self ; Turci_PRL_2021 , and they can potentially lead to a local enhancement in particle displacement and an overall increase in the effective kinetic energy. We first quantify this effect by calculating the effective kinetic energy as a function of distance from the wall, $\langle E_{k}(y)\rangle_{x}=\frac{1}{2}\langle\left(\frac{\Delta{\bf r}(y)}{\Delta t}\right)^{2}\rangle_{x}$ on which the dense layer is formed. Here, $\langle\rangle_{x}$ symbolizes averaging over the particles whose vertical distance from the wet wall lies between $y$ and $y+\delta y$ ($\delta y=1.6\sigma$), and we take $\Delta t=0.1$. This choice of $\Delta t$ ensures that we capture the local dynamics of active particles while they remain within the dense layer. On the other hand, a smaller $\Delta t$ should lead to capturing the effect of translational noise due to velocity fluctuations. In Fig. 4(a), we plot the scaled effective energy $E_{k}(y)/E_{k0}$, where $E_{k0}=v_{0}^{2}/2$, the kinetic energy corresponding to free self-propulsion, as a function of the distance from the liquid-vapor interface, $y-\langle h\rangle$, where $\langle h\rangle$ is the mean position of the liquid-vapor interface (see Appendix G for details). As shown in Fig. 4(a), we do not observe any significant effect of $D_{r}^{w}$ in $\langle E_{k}\rangle_{x}$, especially inside the dense layer, $y<\langle h\rangle$. This indicates that the average kinetic energy of the particles in the dense layer is not significantly affected by an increase in $D_{r}^{w}$. As a next step, we quantify the probability distribution of the kinetic energy, $p(E_{k}/E_{k0})$ for the particles within the dense layer, whose vertical distance from the wet wall is confined within the distance to the liquid-vapor interface (Appendix F). Interestingly, we find a qualitative difference in this distribution between systems with different $D_{r}^{w}$(Fig 4(b)). For large $D_{r}^{w}$, $p(E_{k}/E_{k0})$ shows a small peak at large $E_{k}/E_{k0}$ values, indicating an enhanced particle displacement with $D_{r}^{w}$. Since the qualitative observation indicates that the frequency of bubble formation also increases with $D_{r}^{w}$, it is plausible that the increase in the fraction of particles with high mobility inside the dense region is related to the formation of bubbles.
@@
Figure 4: (a) Normalized kinetic energy of particles $E_{k}(y)/E_{k0}$ as a function of distance from the interface $|y-\langle h\rangle|$ for the case of the wet walls, plotted for increasing values of $D_{r}^{w}$ in the symmetric to asymmetric wetting transition regime at Pe =150. The grey line at $y=0$ indicates the interface location. A slight variation of the energy in the vapor phase is related to the difference in particle densities. (b) Distribution of the normalized kinetic energy showing higher weightage for having high kinetic energy for increasing $D_{r}^{w}$. (Inset) Increase in bubble formation with increasing $D_{r}^{w}$ for the case of the wet walls plotted for the symmetric to asymmetric wetting transition regime. This has been shown here as the minimum in the probability distribution of the local density, focusing on the low density region to track bubbles, plotted for $Pe=150$.

### v2

@@
We investigate the wetting transitions displayed by the collection of active Brownian particles (ABPs) confined within rigid, impenetrable, flat walls. In our computational study using Brownian dynamics simulations, the wall-particle interactions are implemented with a short-range repulsive potential. Our analyses reveal that an enhanced rotational diffusion at the walls can be used as a control parameter for wetting transitions in the dense aggregates of active particles at the wall. Increasing the wall rotational diffusion destabilizes a uniform, complete wetting state, and the aggregate shows morphological transitions. We observe a sequence of morphological transitions with an increase in wall rotational diffusion: symmetric complete wetting (SCW), asymmetric complete wetting (ACW), partial wetting (PW) with droplet formation, and drying. We compute the contact angle in the PW state as a function of activity and rotational noise. Our analysis indicates that these transitions are linked to enhanced kinetic energy fluctuations of particles and bubble formations in the dense state. We further characterize the nature of these transitions by systematically analyzing an order parameter. Our work shows that modifying local reorientation rates alone is sufficient to induce wetting transitions in active systems.
@@ I Introduction
While the properties of passive fluids on solid substrates have been extensively studied, there is a growing interest in exploring the aggregation and spreading of several active and biological systems on solid substrates and membranes Douezan_PNAS_2011 ; Perez-Gonzalez_NatPhys_2019 ; pallares2023stiffness ; Kusumaatmaja_JCB2021 ; Agudo-Canalejo2021 ; Alert_Langmuir_2019 ; Zhao_2021 ; melaugh2016shaping ; ponisch2019bacterial ; fins2024steer ; Sandor_JCP_2017 ; reichhardt2014absorbing . In addition, the spread of bacterial species on substrates has also been studied in this context melaugh2016shaping ; ponisch2019bacterial . Unlike their passive counterparts, such living systems continuously convert energy into mechanical motion and hence far from equilibrium. The collective behavior of such self-motile systems has been theoretically studied using minimal particle-based models. It is well known that such systems display bulk phase separation at sufficiently large density and motility, called Motility-induced phase separation (MIPS) Redner2013 ; Fily2012 ; Cates2015 . They also display wall accumulation under confinement, even in the absence of explicit adhesive interactions Bechinger2016 ; lee2017interface ; Elgeti2013 ; caprini2018active . Several numerical studies have investigated the nature of wall aggregation of self-propelling particles sepulveda2017wetting ; SepulvedaSoto_PRE2018_activewetting ; neta2021wetting ; Turci_PRL_2021 ; Das2020_SM ; Das2020_PRE ; Turci_SM_2024 . It has been shown that the particle aggregates display a continuous transition from a complete wetting to a partial wetting state on rigid, impenetrable walls as the particle reorientation rate is increased sepulveda2017wetting ; SepulvedaSoto_PRE2018_activewetting . Such wetting transitions of active aggregates are also observed when the substrate properties are altered. The effect of wall permeablilty Turci_PRL_2021 ; Turci_SM_2024 and porosity Das2020_SM ; Das2020_PRE , and particle inertia caprini2024dynamical are explored in some detail. Further, a mechanical definition of liquid-solid surface tension has been derived recently zakine2020surface ; zhao2024active . An intriguing morphological transition is observed in the aggregates of particles on the porous walls. With a slight increase in the particle crossing rate, the aggregates display a morphological transition from a uniformly spread state to a liquid droplet with a non-zero contact angle Das2020_SM ; Das2020_PRE , displaying a striking similarity to equilibrium wetting transitions. Apart from systems of self-propelling particles, studies have also been conducted on wetting transitions in other non-equilibrium models Hinrichsen_PRL_1997 ; Hinrichsen_PRE_2003 .
@@ II Model and simulation details
We have conducted the simulations with system sizes varying from $N=4455$ to $N=38400$, while keeping the area fraction $\varphi=N\pi\sigma^{2}/4L_{x}L_{y}$ fixed at $\varphi\simeq 0.3$. At this $\phi$, the system is already in the liquid-gas coexistence region, as shown in the previous studies on ABP models with 2D periodic boundary conditions siebert2018critical ; Bialke2015 ; levis2017active . However, the probability of a spontaneous phase separation in the bulk is still very low at this density even for large Pe Redner2013 ; levis2017active . Under confinement, the phase separation is expected to occur only at the walls. Most of our simulations are conducted for $N=38400$ in a rectangular box with $L_{x}/\sigma=617$ and $L_{y}/\sigma=162$. The simulations were run up to time $T/t_{0}=3100$, for a range of $Pe=0-300$, maximum up to $10$ independent runs. For simulations, we use a time step $10^{-5}t_{0}$ for $Pe<180$ and $5\times 10^{-6}t_{0}$ for $Pe\geq 180$. We vary the rotational diffusion near walls $D_{r}^{w}$ in the range of $0-4335$. The upper limit of $D_{r}^{w}$ is limited such that the time-step of integration is at least an order lesser than the time to rotate for escaping a flat surface $\pi/2D_{r}^{w}$. A summary of the parameters used can be found in Table I.
We have not observed a bulk phase separation for any of the parameter values. However, the dense aggregation on walls shows variety of morphological states as we vary $D_{r}^{w}$ for different Pe.
@@ III Results
Figure 1: (a) Examples of cluster morphology with rigid, impenetrable walls for $N=38400$ and $Pe=240$: (i) symmetric complete wetting ($D_{r}^{w}/D_{r}=10$), (ii) asymmetric complete wetting ($D_{r}^{w}/D_{r}=45$), (iii) partial wetting ($D_{r}^{w}/D_{r}=125$) (b) Two-dimensional steady-state density distribution, with a color bar denoting the local density $\rho(x,y)$ of the states in (a). (c) Vacancy fraction $\ell$ as a function of $D_{r}^{w}$ for five independent runs. The error bars denote the standard deviation over time. The symbols denote symmetric complete wetting (SCW)($\blacksquare$), asymmetric complete wetting (ACW)(), partial wetting (PW)(
@@
When $D_{r}^{w}$is increased, keeping the Pe same, particles no longer aggregate symmetrically on both walls; instead, we observe a uniform aggregation on one of the walls and no aggregation on the opposite wall (see Fig. 1(a)(ii)). In this state, one of the liquid-vapor interfaces formed in SCW is replaced by a wall-vapor interface. Although a simultaneous occurrence of all three interfaces (liquid-vapor, wall-liquid, and wall-vapor) is a feature of a partial wetting state, the wet wall is covered by a uniform, dense layer of particles, unlike what was typically observed in equilibrium partial wetting (see supplementary MOVIE2). We also observe that the thickness of such a dense layer is approximately twice that of the aggregate thickness formed in SCW and the fraction of particles in the dense state is approximately the same. We term this state, asymmetric complete wetting (ACW). Such asymmetric states have previously been observed in simulations of Lennard-Jones fluids nijmeijer1990wetting ; sikkenk1988simulation . We have measured the local density of the dense and dilute regions after the phase separation and compared it with the co-existence densities of the bulk systems (see Fig. S8). While the vapour phase density is close to the bulk co-existence density, the liquid phase density is much higher compared to the bulk systems, as the presence of a rigid wall leads to an enhanced compression of the dense state.
@@
For both ACW and PW, the particle condensation is initiated at different locations on the wall for different independent realizations. Once such a condensed layer is formed on a particular wall, it does not switch between the walls during the entire simulation time. However, when a droplet is formed in the PW case, it can gradually change its location along the wall as evident in MOVIE3. However, we observe a completely different behaviour in the absence of interparticle interaction (Fig. S7). Since interactions are necessary for a phase separation, we do not observe a thick layer of particles on the wall. Therefore, we also do not observe morphological transitions in such cases.
@@ III.1 Density profile and vacancy fraction
Although the instantaneous steady-state configurations clearly indicate a morphological transition of the dense states, we also calculate the time-averaged density distribution $\rho(x,y)$ of ABPs for the characterization of these states (see Supplementary Material for details). As shown in Fig. 1(b)(i)-(iii), $\rho$ provides a clear distinction between liquid and vapor states and relatively smooth liquid-vapor interfaces for all the morphological states. While SCW and ACW states provide a flat liquid-vapor interface, the PW state shows a curved interface.
We further calculate the vacancy fraction $\ell$, which quantifies the fraction of the length of the wall where the dense layer of particles is absent. In other words, $\ell$ quantifies the fractional length of the wall-vapor interfaces (see Supplementary Material for details). The value of $\ell$ is an indicator of the type of morphological state of the dense aggregates. If $\ell\simeq 0$, the system is in SCW state. In the case of ACW, where one of the walls is dry, $\ell\simeq 0.5$. For partial wetting (droplets) $10.5<\ell<$ 1 and for completely dry state $\ell\simeq 1$.
@@ III.2 State diagram
The morphological states of the dense aggregates shown in Fig. 1 are for a fixed activity ($Pe=240$). However, the qualitative behavior of the system is also influenced by the particle activity. In Fig. 2, we summarize the states as a function of $Pe$ and $D_{r}^{w}/D_{r}$. A novel feature observed in the simulation is the formation of the droplet in the PW state, making it crucial to identify the parameter range where partial wetting occurs. Therefore, we introduce a slightly different definition for vacancy fraction, called the wet-wall vacancy fraction $\ell_{w}$, which is calculated only for walls that are partially or completely wet. With this definition, both SCW and ACW are treated as morphologically equivalent states with the same $\ell_{w}$. Thus, ($\ell_{w}\simeq 0$) for both SCW and ACW and for PW state, $\ell_{w}>0$. However, we keep track of ACW and SCW by estimating $\ell$ and summarize this information in Fig. 2.
Figure 2: State diagram showing steady-state morphology of dense clusters with the variation of activity $Pe$ and scaled enhanced rotational diffusion near the wall $D_{r}^{w}/D_{r}$ for the number of active particles, $N=38400$. The color coding in the panel indicates the wet-wall vacancy fraction, $\ell_{w}$ considered for the wet wall cases, except for states showing complete drying for all realizations. The system exhibits symmetric complete wetting $\blacksquare$ (Fig.1(a)(i)), asymmetric complete wetting (Fig.1(a)(ii)), partial wetting (Fig.1(a)(iii))
@@
For $Pe\lesssim 60$, we do not observe any aggregation, and the system is in a completely dry state for all $D_{r}^{w}/D_{r}$. For higher activity ($Pe\gtrsim 100$), we observe SCW states for low $D_{r}^{w}/D_{r}$ ($D_{r}^{w}/D_{r}$ $\lesssim 33.3$) and other aggregated states at higher $D_{r}^{w}/D_{r}$. The droplet formations in PW states are observed for a range of $D_{r}^{w}/D_{r}$ for sufficiently large Pe, which is distinguished by a higher $\ell_{w}$. This range becomes narrower as $Pe$ is decreased until it disappears for $Pe\lesssim 100$. This region is also marked by the existence of bi-stable states, where PW states coexist with ACW states. In the state diagram Fig. 2, we observe the appearance of re-entrant wetting behavior for $D_{r}^{w}/D_{r}\geq 500$.
@@ III.3 Contact angle of droplets
Once we identify the parameter range where the droplet states are observed, we analyze how the droplet morphology is influenced by both $Pe$ and $D_{r}^{w}$. For this, we use the time-averaged density profile, as it clearly identifies the liquid-vapor interface and its shape. First, we note that for a fixed $D_{r}^{w}$, increasing activity leads to a change in droplet shape, as illustrated in Fig. 3(a). For small $Pe$, the droplet is relatively flat upon increasing the $Pe$, and the droplet curvature increases until it saturates at large $Pe$ (Fig.3(a)). We also calculate the contact angle $\theta$ from the local tangent of the liquid-vapor interface at the liquid-wall contact point for different parameters. The contact angle is low for a relatively flat aggregate, and it increases as the interfacial curvature increases (see Fig. S4) In Fig. 3(b), we plot $\theta$ as a function of $D_{r}^{w}/D_{r}$ for a fixed $Pe$. Similar to the trend observed with the change in $Pe$, we find that $\theta$ is sensitive to an increase in $D_{r}^{w}/D_{r}$ for relatively small values of $D_{r}^{w}/D_{r}$, and it saturates around $\pi/6$ at large $D_{r}^{w}/D_{r}$. We note that we never observe $\theta$ approaching $\pi/2$ for any of the simulation parameters.
Figure 3: (a) Time-averaged density profile for droplets keeping $D_{r}^{w}/D_{r}=180$ for increasing $Pe$, showing gradual increase in curvature of droplets and the contact angle. (b) Contact angle $\theta$ (in °$\degree$) for droplets for increasing $D_{r}^{w}/D_{r}$ for $Pe=240$ and $Pe=300$. The dotted lines are guide to the eye.
@@ III.4 Analysis of asymmetric wetting transition
We have observed from the simulations that the dense state often encompasses ‘bubbles’, temporarily formed compact domains with low density (see Fig. S5). Qualitative analyses of simulation results indicate an increase in the number of bubbles formed within the dense layer. Such bubble formations are reported inside the dense regions formed by MIPS and in the presence of walls  martinroca2021mips ; caporusso2020mips ; shi2020self ; Turci_PRL_2021 , and they can potentially lead to a local enhancement in particle displacement and an overall increase in the effective kinetic energy. We first quantify this effect by calculating the effective kinetic energy as a function of distance from the wall, $\langle E_{k}(y)\rangle_{x}=\frac{1}{2}\langle\left(\frac{\Delta{\bf r}(y)}{\Delta t}\right)^{2}\rangle_{x}$ on which the dense layer is formed. Here, $\langle\rangle_{x}$ symbolizes averaging over the particles whose vertical distance from the wet wall lies between $y$ and $y+\delta y$ ($\delta y=1.6\sigma$), and we take $\Delta t=0.1$. This choice of $\Delta t$ ensures that we capture the local dynamics of active particles while they remain within the dense layer. On the other hand, a smaller $\Delta t$ should lead to capturing the effect of translational noise due to velocity fluctuations. In Fig. 4(a), we plot the scaled effective energy $E_{k}(y)/E_{k0}$, where $E_{k0}=v_{0}^{2}/2$, the kinetic energy corresponding to free self-propulsion, as a function of the distance from the liquid-vapor interface, $y-\langle h\rangle$, where $\langle h\rangle$ is the mean position of the liquid-vapor interface (see Supplementary Material). As shown in Fig. 4(a), we do not observe any significant effect of $D_{r}^{w}/D_{r}$ in $\langle E_{k}\rangle_{x}$, especially inside the dense layer, $y<\langle h\rangle$. This indicates that the average kinetic energy of the particles in the dense layer is not significantly affected by an increase in $D_{r}^{w}/D_{r}$. As a next step, we quantify the probability distribution of the kinetic energy, $p(E_{k}/E_{k0})$ for the particles within the dense layer, whose vertical distance from the wet wall is confined within the distance to the liquid-vapor interface (see Supplementary Material). Interestingly, we find a qualitative difference in this distribution between systems with different $D_{r}^{w}/D_{r}$(Fig 4(b)). For large $D_{r}^{w}/D_{r}$, $p(E_{k}/E_{k0})$ shows a small peak at large $E_{k}/E_{k0}$ values, indicating an enhanced particle displacement with $D_{r}^{w}$. Since the qualitative observation indicates that the frequency of bubble formation also increases with $D_{r}^{w}$, it is plausible that the increase in the fraction of particles with high mobility inside the dense region is related to the formation of bubbles.
@@
Figure 4: (a) Normalized kinetic energy of particles $E_{k}(y)/E_{k0}$ as a function of distance from the interface $|y-\langle h\rangle|$ for the case of the wet walls, plotted for increasing values of $D_{r}^{w}/D_{r}$ in the symmetric to asymmetric wetting transition regime at Pe =150. The grey line at $y=0$ indicates the interface location. A slight variation of the energy in the vapor phase is related to the difference in particle densities. (b) Distribution of the normalized kinetic energy showing higher weightage for having high kinetic energy for increasing $D_{r}^{w}/D_{r}$. (Inset) Increase in bubble formation with increasing $D_{r}^{w}/D_{r}$ for the case of the wet walls plotted for the symmetric to asymmetric wetting transition regime. This has been shown here as the minimum in the probability distribution of the local density, focusing on the low density region to track bubbles, plotted for $Pe=150$.